how do i simplify 2+8i/3i
Answer by stanbon(48540)

simplify 2+8i/3i
------
The conjugate of a+bi is a-bi
You have (2+8i)/(0+3i)
Multiply numerator and denominator by (0-3i) to get:
[(2+8i)(-3i)]/[(3i)(-3i)]
=(24-6i)/9
=(8-2i)/3
=(8/3)-(2/3)i
================
Cheers,
Stan H.